# A Survey of 3D Reconstruction with Event Cameras

### v2

@@ V-B Multimodal Methods with Neural Radiance Fields
Qi et al. [152] E2NeRF 2023-10 Blurry RGB, Event stream EAF Self-collected [152]
Ma et al. [139] Deformable Event-NeRF 2023-10 Blurry RGB, Event Stream Event-by-event HS-ERGB [162]
Cannici et al. [154] Ev-DeblurNeRF 2024-06 Blurry RGB, Event stream EAF Ev-DeblurBlender [163]
Qi et al. [155] E3NeRF 2024-08 Blurry RGB, Event stream EGA Self-collected [155]
Tang et al. [156] LSE-NeRF 2024-09 Blurry RGB, Event stream EAF Self-collected [156], EVIMOv2 [164]
Li et al. [160] BeNeRF 2024-10 Blurry RGB, Event stream EAF Qi et al. [152]
Qi et al. [157] EBAD-NeRF 2024-10 Blurry RGB, Event stream EAF Self-collected [157]
Chen et al. [158] Event-ID 2024-10 Blurry RGB, Event stream EAF Self-collected [158]
Rudnev et al. [153] Dynamic EventNeRF 2024-12 Blurry RGB, Event stream EAF Self-collected [153]
Chen et al. [161] EvHDR-NeRF 2025-04 LDR RGB, Event stream EAF HDR-NeRF dataset [165]
Dynamic scene reconstruction: Some methods reconstruct dynamic scenes by combining event streams with deformable or time-conditioned NeRF frameworks. In 2024, Ma et al. [139] proposed DE-NeRF, which is the first deformable NeRF framework that fuses RGB images and event data. It combines continuous pose estimation and a learnable deformation field. The method enables high-quality dynamic scene reconstruction and novel view synthesis. In the same year, Rudnev et al. [153] proposed Dynamic EventNeRF, which uses time-conditioned NeRF models and introduces an event accumulation damping mechanism. The method achieves photorealistic synthesis in low-light and high-speed dynamic scenes using only event cameras and sparse blurry RGB frames.
V-C Multimodal Methods with 3D Gaussian Splatting
Figure 15: Overview of multimodal methods with 3D Gaussian Splatting. (a) The method structure from Deguchi et al. that performs static scene reconstruction [29]. (b) The method structure from Xu et al. that performs dynamic scene reconstruction [166].
TABLE IX: Event cameras in multimodal system: 3DGS-based methods
@@ V-C Multimodal Methods with 3D Gaussian Splatting
Yu et al. [142] EvaGaussians 2024-05 Blurry RGB, Event stream EGA Self-collected [142]
Xiong et al. [167] Event3DGS 2024-06 Blurry RGB, Event stream EGA Self-collected [167]
Weng et al. [55] EaDeblur-GS 2024-07 Blurry RGB, Event stream EGA Qi et al. [152]
Liao et al. [56] EF-3DGS 2024-10 Blurry RGB, Event stream EAF Tanks and Temples [168]
Deguchi et al. [29] E2GS 2024-10 Blurry RGB, Event stream EAF Self-collected [29]
Xu et al. [166] EventBoosted-3DGS 2024-11 Blurry RGB, Event stream EGA Self-collected [166]
Huang et al. [169] Ev3DGS 2024-12 Blurry RGB, Event stream EAF Qi et al. [152]
Wu et al. [170] SweepEvGS 2024-12 Static RGB, Event stream EGA Self-collected [170]
@@
Lee et al. [171] Sensor Fusion Splatting 2025-02 RGB, Event stream, Depths EGA Self-collected [171]
Matta et al. [172] BeSplat 2025-03 Blurry RGB, Event stream EAF BeNeRF dataset [173]
Deng et al. [174] EBAD-Gaussian 2025-04 Blurry RGB, Event stream EAF Qi et al. [157]
Figure 16: Multimodal 3D Gaussian reconstruction based on blurry RGB and event data originates from the E2GS method [29] under CC BY 4.0. @IEEE 2024. The event modality assists in deblurring scene rendering.
@@
Multimodal Event-based Gaussian Splatting methods achieve high-quality 3D reconstruction by combining event streams with RGB images. These multimodal approaches leverage the high temporal resolution of events and the rich appearance information of RGB frames to improve geometric accuracy and colour fidelity. Table IX and Figure 15 provide an overview of these methods. Based on the type of scene being reconstructed, we categorise these methods into the following two types:
@@
Static scene reconstruction: Some methods focus on static scene reconstruction by using event streams to compensate for motion blur and improve 3D reconstruction from blurry images. In 2024, Yu et al. [142] proposed EvaGaussians, which introduces learnable camera pose offsets and jointly optimises blurry image trajectories and 3D Gaussians. It improves reconstruction accuracy under severe motion blur conditions. Later that year, Weng et al. [55] proposed EaDeblur-GS, incorporating an Adaptive Deviation Estimator (ADE) [175, 176] network and blur-specific losses. It enables real-time and high-fidelity 3D reconstruction from extremely blurred inputs. In the same year, Deguchi et al. [29] proposed E2GS, which combines Event-based Double Integral (EDI) [177] and event rendering loss to enhance blurry image recovery and achieve 3D Gaussian reconstruction. Also in 2024, Huang et al. [169] proposed Ev3DGS, which introduces blur and event rendering losses to guide 3D Gaussian Splatting, enabling fast and accurate colour reconstruction from blurry RGB images and event streams. In 2025, Matta et al. [172] proposed BeSplat, which reconstructs sharp and colour-consistent 3D scenes from a single blurry RGB image and event stream by jointly optimising scene representation and camera motion with event-guided spatio-temporal lifting. Later, Deng et al. [174] proposed EBAD-Gaussian, which leverages complementary image and event modalities and introduces event-driven bundle adjustment with motion blur modelling to jointly optimise Gaussians and camera poses, enabling sharp and physically consistent 3D reconstruction under severe motion blur. Figure 16 provides a visualisation of its result compared to RGB-based 3D Gaussian Splatting [69].
@@
A method uses grayscale images as auxiliary input. Compared to other multimodal approaches that combine RGB images, this event-driven strategy is more lightweight. In 2024, Wu et al. [170] proposed SweepEvGS, which is the first 3DGS framework that reconstructs macro- and micro-scale radiance fields from a single sweep of event data. It utilises event-based supervision and structure loss to achieve efficient and robust novel view synthesis.
@@
Dynamic scene reconstruction: Some methods achieve dynamic 3D reconstruction either by modelling continuous time supervision from event streams [167, 56] or by introducing non-rigid deformation fields to explicitly represent scene dynamics [166, 171]. In 2024, Xiong et al. [167] proposed Event3DGS, which introduces sparsity-aware sampling and a progressive training strategy. It reconstructs geometrically consistent and efficient 3D structures from event streams under high-speed egomotion. In the same year, Liao et al. [56] proposed EF-3DGS, which is the first event-aided 3DGS framework that supports free-trajectory rendering in dynamic scenes. It enhances reconstruction accuracy and robustness in dynamic scenes by combining a Linear Event Generation Model (LEGM) [178, 179, 180] with a contrast maximisation-based image sharpening strategy [181, 182, 183, 184]. Also in 2024, Xu et al. [166] proposed Event-Boosted Deformable 3D Gaussians, which is the first deformable 3DGS framework incorporating event cameras. It applies a GS-threshold joint modelling strategy and a dynamic-static decomposition mechanism to achieve high-quality and efficient dynamic reconstruction. In 2025, Lee et al. [171] proposed Sensor Fusion Splatting, which fuses RGB images, event streams, and depth maps from an RGB-D camera with deformable Gaussians and modality-specific losses, enabling high-quality 3D reconstruction in dynamic scenes.
VI Datasets
@@
Although numerous event-based vision datasets have been introduced, only a limited subset is suitable for 3D reconstruction, and many of them are not specially designed for 3D reconstruction. Many recent works rely on private datasets (e.g., [55]) or synthesise events from RGB videos using simulators without releasing the resulting data (e.g., [142], [166], [72]). As a result, only a small number of publicly available datasets provide the dense depth, stereo disparity, or sub-millimetre trajectories required for rigorous evaluation of reconstruction performance. Table X summarises the main publicly available datasets, and Table XI provides example visualisation of sensors, RGB frames, event frames, and labels from the representative datasets. These datasets naturally fall into four categories based on their data acquisition setting, type of geometric supervision, and target reconstruction tasks: (1) outdoor navigation datasets with LiDAR or visual-inertial ground truth, (2) indoor object and human reconstruction datasets with motion capture or laser scans, (3) synthetic datasets with dense annotations, and (4) photometric benchmarks with accurate camera poses but without metric depth.
@@
TABLE X: Publicly available datasets for 3D reconstruction tasks with event cameras. (Abbr.: E = event stream, RGB = image frames, Li = LiDAR, IMU = inertial unit, Vcn = Vicon.)
@@
MVSEC [80] RA-L (2018) Real Stereo 346$\times$260 E, Li, IMU, GPS Point Cloud 30 GB
DHP19 [185] CVPR (2019) Real 4$\times$346$\times$260 E, Vcn 13‑joint skeleton 30 GB
DSEC [93][83] RA-L (2021) Real 2$\times$640$\times$480 E, 2$\times$1.4 MP RGB, Li, IMU Depth Map (LiDAR) 150 GB
TUM‑VIE [124] IROS (2021) Real Stereo 1280$\times$720 E, IMU 200 Hz, Vcn Grayscale Frames 300 GB
ViViD++ [186] RA-L (2022) Real Mono E + Thermal + Li Pose 200 GB
MOEC‑3D [106] ECCV (2022) Real Mono E, laser mesh Mesh 30 GB
EVIMO‑2 [164] Arxiv (2022) Real 3$\times$640$\times$480 E, RGB, 2 IMU, Vcn Object Pose 350 GB
EventScape [187] RA-L (2021) Synthetic CARLA (E+RGB) Dense Depth Maps 70 GB
EventNeRF [35] CVPR (2023) Synthetic Mono E + RGB Refs RGB Frame 18 GB
@@
SynthEVox3D [95] ICVR (2023) Synthetic Mono E (E2V) from Blender Simulator Voxel Grid 32 GB
SEVD [188] Arxiv (2024) Synthetic CARLA multiview E+RGB Depth Map, Masks 300 GB
DAVIS240C [101] IJRR (2017) Real + Synthetic Mono 240$\times$180 E, IMU Pose 10 GB

### v3

@@ 5.1 Multimodal Methods with Traditional Outputs
Qi et al. [[N5]] E2NeRF 2023-10 Blurry RGB, Event stream EAF Self-collected [[N5]]
Ma et al. [[N9]] Deformable Event-NeRF 2023-10 Blurry RGB, Event Stream Event-by-event HS-ERGB [HS-ERGB]
Cannici et al. [[N6]] Ev-DeblurNeRF 2024-06 Blurry RGB, Event stream EAF Ev-DeblurBlender [[N6-D]]
Qi et al. [[N10]] E3NeRF 2024-08 Blurry RGB, Event stream EGA Self-collected [[N10]]
Li et al. [[XN5]] BeNeRF 2024-10 Blurry RGB, Event stream EAF Qi et al. [[N5]]
Qi et al. [[N11]] EBAD-NeRF 2024-10 Blurry RGB, Event stream EAF Self-collected [[N11]]
Chen et al. [[N12]] Event-ID 2024-10 Blurry RGB, Event stream EAF Self-collected [[N12]]
Tang et al. [[N15]] LSE-NeRF 2025-03 Blurry RGB, Event stream EAF Self-collected [[N15]], EVIMOv2 [EVIMO2]
Chen et al. [[XN3]] EvHDR-NeRF 2025-04 LDR RGB, Event stream EAF HDR-NeRF dataset [[XN3-D]]
Rudnev et al. [[N14]] Dynamic EventNeRF 2025-06 Blurry RGB, Event stream EAF Self-collected [[N14]]
5.1.1 Combine with Structured light
Structured light is an active 3D sensing technique that projects coded patterns onto a surface and reconstructs depth via triangulation [StructureLightIntro]. When integrated with event cameras, structured light systems can achieve robust, high-speed depth sensing under challenging conditions. Recent works have introduced several types of encoding and fusion strategies to use both spatial and temporal features of event streams.
In 2015, Matsuda et al. [[XDMT1]] proposed MC3D, one of the earliest event-based structured light systems. The method correlates laser scan timing with event timestamps, achieving high-precision depth reconstruction with per-pixel single-shot efficiency. In 2018, Leroux et al. [[DMT1]] proposed event-based structured light systems by projecting frequency-tagged light patterns onto the scene. Each spatial region is illuminated with a unique modulation frequency, and the event camera decodes depth by associating detected frequencies with pixel locations. In 2021, Huang et al. [[DMT2]] combined structured light projection with digital image correlation (DIC) [[DMT2-23], [DMT2-24]] for high-speed scanning. Also in 2021, Muglikar et al. [[XDMT2]] proposed ESL, which estimates depth by maximising spatio-temporal consistency between a laser projector and an event camera. By processing events in local space-time regions, their method improves robustness to noise. ESL accurately estimates depth in challenging scenes but is not real-time due to high computational cost. In 2023, Xiao et al. [[DMT4]] employed alternating binary speckle patterns and DIC-based stereo matching for fast and accurate reconstruction. Also in 2023, Fu et al. [[DMT5]] introduced spatio-temporal coding (STC) with an enhanced matching scheme for improved stereo robustness. In 2024, Li et al. [[DMT6]] proposed eFPSL, using time-frequency analysis to extract high-SNR fringe maps from events and an event-count-based shadow mask to reduce errors.
5.1.2 Combine with RGB-D
Some methods fuse event data with RGB-D sensors for improved scene understanding from depth and RGB information. In 2014, Weikersdorfer et al. [[XDMT3]] proposed EB-SLAM-3D, a novel event-based 3D SLAM algorithm using a D-eDVS, enabling low-power, low-latency mapping with a sparse voxel grid at 20× real-time speed. In 2022, Zuo et al. [[DMT3]] proposed DEVO, combining time surface maps from events and depth supervision from a calibrated sensor. Their system performs semi-dense 3D-2D edge alignment to estimate poses and incrementally build point clouds under fast motion and poor lighting.
5.2 Multimodal Methods with Neural Radiance Fields
Figure 15: Overview of multimodal methods with 3D Gaussian Splatting. (a) The method structure from Deguchi et al. that performs static scene reconstruction. (b) The method structure from Xu et al. that performs dynamic scene reconstruction.
Table 9: Event cameras in multimodal system: 3DGS-based methods
@@ 5.2 Multimodal Methods with Neural Radiance Fields
Yu et al. [[G2]] EvaGaussians 2024-05 Blurry RGB, Event stream EGA Self-collected [[G2]]
Xiong et al. [[G4]] Event3DGS 2024-06 Blurry RGB, Event stream EGA Self-collected [[G4]]
Weng et al. [[G3]] EaDeblur-GS 2024-07 Blurry RGB, Event stream EGA Qi et al. [[N5]]
Liao et al. [[G13]] EF-3DGS 2024-10 Blurry RGB, Event stream EAF Tanks and Temples [Tanks]
Deguchi et al. [[G1]] E2GS 2024-10 Blurry RGB, Event stream EAF Self-collected [[G1]]
Xu et al. [[G15]] EventBoosted-3DGS 2024-11 Blurry RGB, Event stream EGA Self-collected [[G15]]
Huang et al. [[XG1]] Ev3DGS 2024-12 Blurry RGB, Event stream EAF Qi et al. [[N5]]
Wu et al. [[G14]] SweepEvGS 2024-12 Static RGB, Event stream EGA Self-collected [[G14]]
@@
Lee et al. [[XG3]] Sensor Fusion Splatting 2025-02 RGB, Event stream, Depths EGA Self-collected [[XG3]]
Matta et al. [[XG4]] BeSplat 2025-03 Blurry RGB, Event stream EAF BeNeRF dataset [[XG4-D]]
Deng et al. [[XG5]] EBAD-Gaussian 2025-04 Blurry RGB, Event stream EAF Qi et al. [[N11]]
Lee et al. [[XG6]] DiET-GS 2025-06 Blurry RGB, Event stream EGA Cannici et al. [[N6]]
@@
Multimodal Event-based NeRF methods combine event streams with RGB images to achieve high-quality coloured 3D reconstruction, leveraging the temporal precision of events and the rich appearance information of RGB frames to enhance structural accuracy and visual fidelity under motion blur and low-light conditions. Table 8 and Figure 14 provide an overview of these methods. Based on the type of scene being reconstructed, we categorise these methods into the following two types:
5.2.1 Static scene reconstruction
Some methods focus on static scene reconstruction by integrating event supervision with blurry RGB images [[N5], [N6], [N10], [N15]], or by designing physically inspired mechanisms [[N11], [N12]] to improve NeRF training stability and output sharpness. In 2023, Qi et al. [[N5]] proposed E²NeRF, which introduces a blur rendering loss and an event rendering loss to improve NeRF training under motion blur and low-light settings. In 2024, Cannici et al. [[N6]] proposed Ev-DeblurNeRF, which uses the Event Double Integral (EDI) [[N6-29]] and a learnable event camera response function (eCRF) to reconstruct sharp NeRFs. It performs well under severe motion blur. Later that year, Qi et al. [[N10]] proposed E3NeRF, which combines blurry images and events with spatial-temporal attention [guo2022attention]. It introduces an event-enhanced rendering loss to guide learning and improves robustness in non-uniform motion and low-light scenes. Later, Li et al. [[XN5]] proposed BeNeRF, which jointly reconstructs photorealistic 3D scenes and estimates camera motion from a single blurry RGB image and event stream, using an event rendering loss and B-spline trajectory representation for end-to-end optimisation. Around the same time, Qi et al. [[N11]] proposed EBAD-NeRF, which introduces event-driven bundle adjustment. It jointly optimises NeRF and camera poses using an intensity-change event loss and a photometric blur loss. Also in 2024, Chen et al. [[N12]] proposed Event-ID, which is the first event-driven framework for intrinsic decomposition. It combines an event-based reflectance model and a multi-view strategy to recover geometry, materials, and lighting under extreme conditions. In 2025, Tang et al. [[N15]] proposed LSE-NeRF, which models sensor response differences with per-time embeddings and event-based reflectance mapping to recover high-quality NeRFs without strict alignment. Later, Chen et al. [[XN3]] proposed EvHDR-NeRF, which models a radiance-based relationship that accounts for exposure time and the camera response function (CRF), enabling HDR 3D reconstruction from single-exposure LDR images and event streams.
5.2.2 Dynamic scene reconstruction
Some methods reconstruct dynamic scenes by combining event streams with deformable or time-conditioned NeRF frameworks. In 2024, Ma et al. [[N9]] proposed DE-NeRF, which is the first deformable NeRF framework that fuses RGB images and event data. It combines continuous pose estimation and a learnable deformation field. The method enables high-quality dynamic scene reconstruction and novel view synthesis. In 2025, Rudnev et al. [[N14]] proposed Dynamic EventNeRF, which uses time-conditioned NeRF models and introduces an event accumulation damping mechanism. The method achieves photorealistic synthesis in low-light and high-speed dynamic scenes using only event cameras and sparse blurry RGB frames.
5.3 Multimodal Methods with 3D Gaussian Splatting
Figure 16: Multimodal 3D Gaussian reconstruction based on blurry RGB and event data originates from the E2GS method under $C$ BY 4.0. @IEEE 2024. The event modality assists in deblurring scene rendering.
Multimodal Event-based Gaussian Splatting methods achieve high-quality 3D reconstruction by combining event streams with RGB images. These multimodal approaches leverage the high temporal resolution of events and the rich appearance information of RGB frames to improve geometric accuracy and colour fidelity. Table 9 and Figure 15 provide an overview of these methods. Based on the type of scene being reconstructed, we categorise these methods into the following two types:
5.3.1 Static scene reconstruction
@@
Some methods focus on static scene reconstruction by using event streams to compensate for motion blur and improve 3D reconstruction from blurry images. In 2024, Yu et al. [[G2]] proposed EvaGaussians, which introduces learnable camera pose offsets and jointly optimises blurry image trajectories and 3D Gaussians. It improves reconstruction accuracy under severe motion blur conditions. Later that year, Weng et al. [[G3]] proposed EaDeblur-GS, incorporating an Adaptive Deviation Estimator (ADE) [[G3-9], [G3-2]] network and blur-specific losses. It enables real-time and high-fidelity 3D reconstruction from extremely blurred inputs. In the same year, Deguchi et al. [[G1]] proposed E2GS, which combines Event-based Double Integral (EDI) [[G1-19]] and event rendering loss to enhance blurry image recovery and achieve 3D Gaussian reconstruction. Also in 2024, Huang et al. [[XG1]] proposed Ev3DGS, which introduces blur and event rendering losses to guide 3D Gaussian Splatting, enabling fast and accurate colour reconstruction from blurry RGB images and event streams. In 2025, Matta et al. [[XG4]] proposed BeSplat, which reconstructs sharp and colour-consistent 3D scenes from a single blurry RGB image and event stream by jointly optimising scene representation and camera motion with event-guided spatio-temporal lifting. Later, Deng et al. [[XG5]] proposed EBAD-Gaussian, which leverages complementary image and event modalities and introduces event-driven bundle adjustment with motion blur modelling to jointly optimise Gaussians and camera poses, enabling sharp and physically consistent 3D reconstruction under severe motion blur. In the same year, Lee et al. [[XG6]] proposed DIET-GS, which also integrates the EDI prior with a pre-trained diffusion model. It employs a two-stage training strategy to constrain 3D Gaussians with event data for accurate colour recovery, while leveraging the diffusion prior to further refine fine-grained details and enhance edge sharpness. Figure 16 provides a visualisation of its result compared to RGB-based 3D Gaussian Splatting [[Gaussian]].
@@
A method uses grayscale images as auxiliary input. Compared to other multimodal approaches that combine RGB images, this event-driven strategy is more lightweight. In 2024, Wu et al. [[G14]] proposed SweepEvGS, which is the first 3DGS framework that reconstructs macro- and micro-scale radiance fields from a single sweep of event data. It utilises event-based supervision and structure loss to achieve efficient and robust novel view synthesis.
5.3.2 Dynamic scene reconstruction
@@
Some methods achieve dynamic 3D reconstruction either by modelling continuous time supervision from event streams [[G4], [G13]] or by introducing non-rigid deformation fields to explicitly represent scene dynamics [[G15], [XG3]]. In 2024, Xiong et al. [[G4]] proposed Event3DGS, which introduces sparsity-aware sampling and a progressive training strategy. It reconstructs geometrically consistent and efficient 3D structures from event streams under high-speed egomotion. In the same year, Liao et al. [[G13]] proposed EF-3DGS, which is the first event-aided 3DGS framework that supports free-trajectory rendering in dynamic scenes. It enhances reconstruction accuracy and robustness in dynamic scenes by combining a Linear Event Generation Model (LEGM) [[G13-10], [G13-13], [G13-56]] with a contrast maximisation-based image sharpening strategy [[G13-11], [G13-15], [G13-34], [G13-38]]. Also in 2024, Xu et al. [[G15]] proposed Event-Boosted Deformable 3D Gaussians, which is the first deformable 3DGS framework incorporating event cameras. It applies a GS-threshold joint modelling strategy and a dynamic-static decomposition mechanism to achieve high-quality and efficient dynamic reconstruction. In 2025, Lee et al. [[XG3]] proposed Sensor Fusion Splatting, which fuses RGB images, event streams, and depth maps from an RGB-D camera with deformable Gaussians and modality-specific losses, enabling high-quality 3D reconstruction in dynamic scenes.
6 Datasets
@@
Although numerous event-based vision datasets have been introduced, only a limited subset is suitable for 3D reconstruction, and many of them are not specially designed for 3D reconstruction. Many recent works rely on private datasets (e.g., [[G3]]) or synthesise events from RGB videos using simulators without releasing the resulting data (e.g., [[G2]], [[G15]], [[S5]]). As a result, only a small number of publicly available datasets provide the dense depth, stereo disparity, or sub-millimetre trajectories required for rigorous evaluation of reconstruction performance. Table 10 summarises the main publicly available datasets. Table 11 provides example visualisation of sensors, RGB frames, event frames, and labels from the representative datasets. In Table 11, all the synthetic datasets are created by Blender [blender], and parts of the figures in the table are referenced from Ghosh et al. with permission [ghosh2024event]).
These datasets naturally fall into four categories based on their data acquisition setting, type of geometric supervision, and target reconstruction tasks: (1) outdoor navigation datasets with LiDAR or visual-inertial ground truth, (2) indoor object and human reconstruction datasets with motion capture or laser scans, (3) synthetic datasets with dense annotations, and (4) photometric benchmarks with accurate camera poses but without metric depth.
@@
Table 10: Publicly available datasets for 3D reconstruction tasks with event cameras. (Abbr.: E = event stream, RGB = image frames, Li = LiDAR, IMU = inertial unit, Vcn = Vicon.)
@@
[…]
ViViD++ [ViViD++] RA-L (2022) Real Mono E + Thermal + Li Pose 200 GB
MOEC‑3D [[MD5]] ECCV (2022) Real Mono E, laser mesh Mesh 30 GB
[…]
EventScape [EventScape] RA-L (2021) Synthetic CARLA (E+RGB) Dense Depth Maps 70 GB
EventNeRF [[N3]] CVPR (2023) Synthetic Mono E + RGB Refs RGB Frame 18 GB
@@
SynthEVox3D [[Dense]] ICVR (2023) Synthetic Mono E (E2V) from Blender Simulator Voxel Grid 32 GB
SEVD [SEVD] Arxiv (2024) Synthetic CARLA multiview E+RGB Depth Map, Masks 300 GB
[…]
6.1 Outdoor navigation datasets